Abstract. We use techniques of Bannai and Sloane to give a new proof that there is a unique (22, 891, 1/4) spherical code; this result is implicit in a recent paper by Cuypers. We also correct a minor error in the uniqueness proof given by Bannai and Sloane for the (23, 4600, 1/3) spherical code.